# Dynamic Color Assignment for Hierarchical Data

### v3

@@ Introduction
Assigning colors to samples according to their class labels and spatial distribution is a common practice in data analysis [24, 29]. A high-quality color assignment should be discriminable and harmonic to ensure clarity and attractiveness in visualizations [16, 28]. However, as the number of classes increases, it is challenging to select a large number of colors that are easily distinguishable from each other but also harmonize together  [49, 13]. A practical solution is to organize classes into a hierarchical structure and then dynamically assign consistent colors across hierarchical levels. This solution not only enhances scalability by reducing the requirement for a large number of distinct colors in a static visualization, but also alleviates cognitive load during data exploration. However, existing color assignment methods do not fully support this solution. Some methods generate color assignment results independently for each hierarchical level, which cannot maintain color consistency across levels [29, 13]. This will disrupt the user’s mental map during exploration. Other methods apply strict color constraints to achieve consistency but sacrifice discriminability and harmony at each level [48, 15]. These limitations highlight the need for a dynamic color assignment method that is capable of maintaining color discriminability and harmony at each level and ensures consistency across different levels during exploration.
@@
  • An open-source implementation of the proposed color assignment method in both C++ and JavaScript, available at https://github.com/Dynamic-Color/Dynamic-Color.
@@ 2 Requirement Analysis
Consider spatial distribution. Our experts also highlighted the importance of considering spatial distribution in generating color assignment results. Taking spatial distribution into account not only enhances data analysis [24, 55] but also improves the aesthetic appeal of the results [39, 25]. E3 and E6 emphasized that the assigned colors of two spatially adjacent classes significantly affect perception and thus deserve careful consideration. For example, enhancing the color discriminability between spatially adjacent classes can make the class boundaries clearer and aid in identifying different classes. Moreover, maintaining color harmony between adjacent classes can produce more visually pleasing results. E5 also noted that assigning similar colors to similar classes would facilitate data understanding and exploration.
Align with hierarchical structures. Four experts also acknowledged that the generated color assignment results should accurately reflect the hierarchical structures within datasets. E1 said, “It is common practice to use similar colors to encode a parent class and its corresponding child classes, which facilitates the identification of hierarchical relationships and keeps the user’s mental map throughout the zooming process.” E4 further emphasized that to avoid misinterpretation of hierarchical relationships, the color differences between child classes of the same parent class should be smaller than those between child classes of different parent classes.
@@ 3.2.2 Color Harmony
The color harmony is then defined as the sum of hue harmony and chroma-lightness harmony: $E_{\mathrm{H}}=E_{\mathrm{Hue}}+E_{\mathrm{LC}}$, which gives satisfactory results in practice.
@@ 3.2.3 Spatial Distribution
Here, $X$ is the set of all samples, and $\Omega_{x_{i}}$ is the set of neighboring samples of sample $x_{i}$ in visualization. We consider three representative types of visualization: scatterplots (point-based), parallel coordinates (line-based), and grid visualizations (area-based). For grid visualizations, we consider eight surrounding cells of the center cell.For scatterplots and parallel coordinates, we use eight nearest neighbors, which is consistent with grid visualizations.The score of each sample pair is $f(x_{i},x_{j})/d(x_{i},x_{j})$. Here, $d(x_{i},x_{j})$ is the spatial distance between samples $x_{i}$ and $x_{j}$ in the visualization. Therefore, a closer pair would have more impact on the objective. $f(x_{i},x_{j})$ is the optimization objective for each pair. Based on the requirement analysis, we consider two modes for this term: difference mode and similarity mode.
The difference mode favors a color pair with a larger perceptual difference between adjacent classes. Therefore, $f(x_{i},x_{j})$ is set as $D(c(x_{i}),c(x_{j}))+P(c(x_{i}),c(x_{j}))$. The first term is the perceptual difference between the corresponding colors of $x_{i}$ and $x_{j}$. The second term measures the pair harmony between colors of $x_{i}$ and $x_{j}$. We use the formula introduced in Ou’s recent work about color pair harmony [41] to calculate this term.
@@ 3.2.4 Optimization
A straightforward way to solve the optimization problem defined in Eq. (1) is simulated annealing. It is chosen because of its flexibility to accommodate multiple objectives and its effectiveness in escaping local optima during the optimization process. First, we use the blue noise sampling [8] to generate the initial color assignment result within the default color range. This technique ensures that colors are evenly distributed, which provides basic discriminability. At each iteration, the algorithm adjusts the color assignment result and re-evaluates the objective value. Adjustments that improve the objective value are always accepted, while those that reduce the objective value are accepted with a progressively decreasing probability over time. However, the simulated annealing algorithm suffers from slow convergence due to the low acceptance rate that comes with the original highly non-convex problem. To accelerate convergence, we combine it with the continuation method. Starting from solely optimizing discriminability, it sequentially incorporates harmony and spatial distribution. Each solution to the previous problem serves as a starting point for the optimization of the subsequent problem. By guiding the optimization process towards more promising regions in the solution space, this method achieves higher acceptance rates and faster convergence [2]. The experimental result shows that our method can generate color assignment results for 30 classes in 1 second, which well supports real-time interaction for users navigating through hierarchical visualizations (see the supplemental material for more details). When incorporating a new optimization goal in each stage of the continuation method, we dynamically set its weighting parameter using loss-balanced task weighting [27]. The basic idea is to ensure that different goals are optimized in a similar progress. Thus, the weighting parameter of the goal with less progress will be increased so that it can be further improved in subsequent iterations. Specifically, in each stage, the weighting parameter of the newly incorporated goal is set as the ratio between the current objective value and the possibly maximal value. After the algorithm converges, the weighting parameter will be fixed, and the continuation method will move to the next stage. Fig. 4 shows the incremental refinement of the color assignment results through each phase of the continuation method. The idea of the continuation method also aligns with the typical process of hand-crafted color assignment design. Initially, users select a set of distinguishable colors tailored to the number of classes. Next, they adjust the hue, chroma, and lightness to improve harmony. Finally, they assign these colors to different classes in the visualization and make slight adjustments to refine the overall visual effects.
@@ 3.3.2 Dynamic Color Range
First, we ensure that the gap must exceed the radii of both spheres, i.e. , $d_{ij}-r_{i}-r_{j}>1\max(r_{i},r_{j})$. However, during the development of our methods, the experts pointed out that the hue plays a more important role in the identification of parent-child relationships, and only considering perceptual differences can sometimes result in misunderstanding. For example, in the left image, the colors are sampled within the ranges of three spheres: green (top), blue (bottom-left), and orange (bottom-right). Although the color on the top-right corner (  ) is closer to the green color than the orange and the blue, there is still a noticeable difference in its hue compared to the green color. This potentially misleads users to perceive it as belonging to a separate parent class, e.g. , a parent class with yellow color. To avoid such misunderstanding, we add an additional restriction on hue in a similar way, which ensures that the gap between two hue ranges must exceed the length of both ranges. Second, we study how to properly determine the radius of the sphere based on the number of child classes. We conduct an experiment to estimate the relationship between the radius and the number of child classes through blue noise sampling [8]. Specifically, given a sphere with a radius $r$, we employ the widely adopted dart-throwing method [8] to sample colors within it until no more discernible colors can be sampled. Here, a discernible color means that the perceptual differences between it and those sampled colors exceed a threshold of 10, which is consistent with the threshold we used in Sec. 3.2.1. Our experimental results indicate a roughly linear relationship between the maximal number of possible colors $n$ and the square of the radius: $n\propto r^{2}$. Therefore, we introduce constraints that $r_{i}/r_{j}=\sqrt{n_{i}}/\sqrt{n_{j}}$, where $r_{i}$ and $r_{j}$ denote the radii of two spheres associated with two parent classes, and $n_{i}$ and $n_{j}$ denote the number of child classes within those spheres. Finally, the radii of the spheres are determined as the maximal radii that adhere to both $d_{12}-r_{1}-r_{2}>\max(r_{1},r_{2})$ and $r_{1}/r_{2}=\sqrt{n_{1}}/\sqrt{n_{2}}$.
@@ 4.1 Quantitative Evaluation
Datasets. We evaluated the quality of different color assignment results across 12 datasets, which have been widely used in recent visualization research [57, 64]. Six of them (MNIST[23], Animals[10], Indian Food[35], Isolet[14], Texture[1], Clothing[59]) are flat datasets with a moderate number of classes, ranging from 10 to 26. These datasets are used to evaluate different flat color assignment methods. The remaining six datasets (Food101 [3], Flowers102 [37], Stanford Cars [21], Caltech256 [17], NABirds [51], ImageNet1k [12]) have a larger number of classes, ranging from 101 to 1000. These datasets are used to evaluate different hierarchical color assignment methods. For datasets with a pre-existing class hierarchy, such as ImageNet, we directly used their hierarchies. For the other datasets, we applied the commonly used hierarchical k-means method [58] to build the class hierarchy in a top-down manner. More details of these datasets are given in the supplemental material.
We used these datasets to create three types of visualization: scatterplots (point-based) [57], parallel coordinates (line-based) [15], and grid visualizations (area-based) [64], which cover the primary types of visualization techniques for data analysis. We also evaluated the generated color assignment results on their own, i.e. , considering the color palettes solely without integrating them into specific visualizations. When evaluating the flat color assignment methods, we assigned colors to all classes. When evaluating hierarchical color assignment methods, we simulated how users explore the hierarchy by randomly choosing a subtree for expansion.
@@ 6 Conclusion
All supplemental materials are available on OSF at https://osf.io/e4b5u/?view_only=68cc67c194c443b498bd2545ef551faa, released under a CC BY 4.0 license. In particular, they include (1) dataset information, (2) the running time of our color assignment methods, (3) additional color assignment results, (4) additional quantitative evaluation results, (5) additional user study results, and (6) the video of our work.

### v5

@@ Introduction
Assigning colors to samples according to their class labels and spatial distribution is a common practice in data analysis [72, 70, 35]. A high-quality color assignment should be discriminable and harmonic to ensure clarity and attractiveness in visualizations [20, 34]. However, as the number of classes increases, it is challenging to select a large number of colors that are easily distinguishable from each other but also harmonize together  [57, 17]. A practical solution is to organize classes into a hierarchical structure and then dynamically assign consistent colors across hierarchical levels. This solution not only enhances scalability by reducing the requirement for a large number of distinct colors in a static visualization, but also alleviates cognitive load during data exploration. However, existing color assignment methods do not fully support this solution. Some methods generate color assignment results independently for each hierarchical level, which cannot maintain color consistency across levels [36, 17]. This will disrupt the user’s mental map during exploration. Other methods apply strict color constraints to achieve consistency but sacrifice discriminability and harmony at each level [56, 19]. These limitations highlight the need for a dynamic color assignment method that is capable of maintaining color discriminability and harmony at each level and ensures consistency across different levels during exploration.
@@
  • An open-source implementation of the proposed color assignment method in both C++ and JavaScript, available at https://github.com/thu-vis/Dynamic-Color.
@@ 2 Requirement Analysis
Consider spatial distribution. Our experts also highlighted the importance of considering spatial distribution in generating color assignment results. Taking spatial distribution into account not only enhances data analysis [29, 63, 67] but also improves the aesthetic appeal of the results [47, 30, 24]. E3 and E6 emphasized that the assigned colors of two spatially adjacent classes significantly affect perception and thus deserve careful consideration. For example, enhancing the color discriminability between spatially adjacent classes can make the class boundaries clearer and aid in identifying different classes. Moreover, maintaining color harmony between adjacent classes can produce more visually pleasing results. E5 also noted that assigning similar colors to similar classes would facilitate data understanding and exploration.
Align with hierarchical structures. As rich hierarchies are ubiquitous in datasets [71, 6, 38], four experts also acknowledged that the generated color assignment results should accurately reflect the hierarchical structures. E1 said, “It is common practice to use similar colors to encode a parent class and its corresponding child classes, which facilitates the identification of hierarchical relationships and keeps the user’s mental map throughout the zooming process.” E4 further emphasized that to avoid misinterpretation of hierarchical relationships, the color differences between child classes of the same parent class should be smaller than those between child classes of different parent classes.
@@ 3.2.2 Color Harmony
The color harmony is then defined as the sum of hue harmony and chroma-lightness harmony, which gives satisfactory results in practice.
@@ 3.2.3 Spatial Distribution
Here, $X$ is the set of all samples, and $\Omega_{x_{i}}$ is the set of neighboring samples of sample $x_{i}$ in visualization. We consider three representative types of visualization: scatterplots [69] (point-based), parallel coordinates [36] (line-based), and grid visualizations [32, 7] (area-based). For grid visualizations, we consider eight surrounding cells of the center cell.For scatterplots and parallel coordinates, we use eight nearest neighbors, which is consistent with grid visualizations.The score of each sample pair is $f(x_{i},x_{j})/d(x_{i},x_{j})$. Here, $d(x_{i},x_{j})$ is the spatial distance between samples $x_{i}$ and $x_{j}$ in the visualization. Therefore, a closer pair would have more impact on the objective. $f(x_{i},x_{j})$ is the optimization objective for each pair. Based on the requirement analysis, we consider two modes for this term: difference mode and similarity mode.
The difference mode favors a color pair with a larger perceptual difference between adjacent classes. Therefore, $f(x_{i},x_{j})$ is set as $D(c(x_{i}),c(x_{j}))+P(c(x_{i}),c(x_{j}))$. The first term is the perceptual difference between the colors of $x_{i}$ and $x_{j}$. The second term measures the pair harmony between colors of $x_{i}$ and $x_{j}$. We use the formula introduced in Ou’s recent work about color pair harmony [49] to calculate this term.
@@ 3.2.4 Optimization
A straightforward way to solve the optimization problem defined in Eq. (1) is simulated annealing. It is chosen because of its flexibility to accommodate multiple objectives and its effectiveness in escaping local optima during the optimization process. First, we use the blue noise sampling [12] to generate the initial color assignment result within the default color range. This technique ensures that colors are evenly distributed, which provides basic discriminability. At each iteration, the algorithm adjusts the color assignment result and re-evaluates the objective value. Adjustments that improve the objective value are always accepted, while those that reduce the objective value are accepted with a progressively decreasing probability over time. However, the simulated annealing algorithm suffers from slow convergence due to the low acceptance rate that comes with the original highly non-convex problem. To accelerate convergence, we combine it with the continuation method. Starting from solely optimizing discriminability, it sequentially incorporates harmony and spatial distribution. Each solution to the previous problem serves as a starting point for the optimization of the subsequent problem. By guiding the optimization process towards more promising regions in the solution space, this method achieves higher acceptance rates and faster convergence [2]. The experimental result shows that our method can generate color assignment results for 30 classes in 1 second, which well supports real-time interaction for users navigating through hierarchical visualizations (see the supplemental material for more details). When incorporating a new optimization goal in each stage of the continuation method, we dynamically set its weighting parameter using loss-balanced task weighting [33, 68]. The basic idea is to ensure that different goals are optimized in a similar progress. Thus, the weighting parameter of the goal with less progress will be increased so that it can be further improved in subsequent iterations. Specifically, in each stage, the weighting parameter of the newly incorporated goal is set as the ratio between the current objective value and the possibly maximal value. After the algorithm converges, the weighting parameter will be fixed, and the continuation method will move to the next stage. Fig. 4 shows the incremental refinement of the color assignment results through each phase of the continuation method. The idea of the continuation method also aligns with the typical process of hand-crafted color assignment design. Initially, users select a set of distinguishable colors tailored to the number of classes. Next, they adjust the hue, chroma, and lightness to improve harmony. Finally, they assign these colors to different classes in the visualization and make slight adjustments to refine the overall visual effects.
@@ 3.3.2 Dynamic Color Range
First, we ensure that the gap must exceed the radii of both spheres, i.e. , $d_{ij}-r_{i}-r_{j}>1\max(r_{i},r_{j})$. However, during the development of our methods, the experts pointed out that the hue plays a more important role in the identification of parent-child relationships, and only considering perceptual differences can sometimes result in misunderstanding. For example, in the left image, the colors are sampled within the ranges of three spheres: green (top), blue (bottom-left), and orange (bottom-right). Although the color on the top-right corner (  ) is closer to the green color than the orange and the blue, there is still a noticeable difference in its hue compared to the green color. This potentially misleads users to perceive it as belonging to a separate parent class, e.g. , a parent class with yellow color. To avoid such misunderstanding, we add an additional restriction on hue in a similar way, which ensures that the gap between two hue ranges must exceed the length of both ranges. Second, we study how to properly determine the radius of the sphere based on the number of child classes. We conduct an experiment to estimate the relationship between the radius and the number of child classes through blue noise sampling [12, 74]. Specifically, given a sphere with a radius $r$, we employ the widely adopted dart-throwing method [12] to sample colors within it until no more discernible colors can be sampled. Here, a discernible color means that the perceptual differences between it and those sampled colors exceed a threshold of 10, which is consistent with the threshold we used in Sec. 3.2.1. Our experimental results indicate a roughly linear relationship between the maximal number of possible colors $n$ and the square of the radius: $n\propto r^{2}$. Therefore, we introduce constraints that $r_{i}/r_{j}=\sqrt{n_{i}}/\sqrt{n_{j}}$, where $r_{i}$ and $r_{j}$ denote the radii of two spheres associated with two parent classes, and $n_{i}$ and $n_{j}$ denote the number of child classes within those spheres. Finally, the radii of the spheres are determined as the maximal radii that adhere to both $d_{12}-r_{1}-r_{2}>\max(r_{1},r_{2})$ and $r_{1}/r_{2}=\sqrt{n_{1}}/\sqrt{n_{2}}$.
@@ 4.1 Quantitative Evaluation
Datasets. We evaluated the quality of different color assignment results across 12 datasets, which have been widely used in recent visualization research [65, 77]. Six of them (MNIST[28], Animals[14], Indian Food[43], Isolet[18], Texture[1], Clothing[65]) are flat datasets with a moderate number of classes, ranging from 10 to 26. These datasets are used to evaluate different flat color assignment methods. The remaining six datasets (Food101 [3], Flowers102 [45], Stanford Cars [26], Caltech256 [21], NABirds [59], ImageNet1k [16]) have a larger number of classes, ranging from 101 to 1000. These datasets are used to evaluate different hierarchical color assignment methods. For datasets with a pre-existing class hierarchy, such as ImageNet, we directly used their hierarchies. For the other datasets, we applied the commonly used hierarchical k-means method to build the class hierarchy in a top-down manner. More details of these datasets are given in the supplemental material.
We used these datasets to create three types of visualization: scatterplots (point-based) [65, 8], parallel coordinates (line-based) [19], and grid visualizations (area-based) [77, 9], which cover the primary types of visualization techniques for data analysis. We also evaluated the generated color assignment results on their own, i.e. , considering the color palettes solely without integrating them into specific visualizations. When evaluating the flat color assignment methods, we assigned colors to all classes. When evaluating hierarchical color assignment methods, we simulated how users explore the hierarchy by randomly choosing a subtree for expansion.
@@ 6 Conclusion
All supplemental materials are available on OSF at https://osf.io/e4b5u/?view_only=68cc67c194c443b498bd2545ef551faa, released under a CC BY 4.0 license. In particular, they include (1) dataset information, (2) the running time of our color assignment methods, (3) additional color assignment results, (4) additional quantitative evaluation results, (5) additional user study results, and (6) the video.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China under grants U21A20469, 61936002, grants from the Institute Guo Qiang, THUIBCS, and BLBCI. The authors would like to thank Jun Yuan, Zhen Li, and Duan Li for their valuable contributions to the discussions, and Yiwei Hou for her assistance in voicing our video.